OpenStudy (anonymous):

Solve for x: 3(x+1)=-2(x-1)+6

OpenStudy (akashdeepdeb):

Use the distributive property!! :D Open the Parenthesis 3(x+1)=-2(x-1)+6 3x + 3 = -2x + 2+6 3x + 2x + 3 -3 = -2x + 2x -3 + 2 + 6 5x = 5 x = 1 Understood? :)

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

:)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."